A collector is using a magnifying lens with a radius of curvature of 21.0 cm to examine the finer details of a butterfly with a magnification of 5.10. At what distance between the butterfly and the lens does she achieve the desired magnification?
Answer #1

SOIOR = 21 cm =절에 f = R 2 m511 - V o 는 - 510 V=-5.10 0 the lone equation using 1 -1 - & ~ 51100 영름 금 (14) = 3 U = -12-5 6 cm